OpenStudy (osanseviero):

Which is the limit of 0.4, 0.44, 0.444, 0.4444...

OpenStudy (anonymous):

Are you talking about an infinite number?

OpenStudy (osanseviero):

I will continue adding and adding and adding...so...does it go to infinite? or is there a limit?

OpenStudy (anonymous):

No there is no limit.

OpenStudy (kira_yamato):

Original SEQ: 0.4, 0.44, 0.444, 0.4444 Difference 1: 0.04 , 0.004 , 0.0004, which is a geometric progression ^^

OpenStudy (kira_yamato):

The difference between the terms has a limit. Just add that to 0.4

OpenStudy (anonymous):

http://openstudy.com/study#/updates/50602f40e4b0583d5cd23d35 this might help you

OpenStudy (anonymous):

1/3 = 0.333... (the 3 repeats forever) 1/7 = 0.14285714285... ( the "142857" repeats forever) 77/600 = 0.128333... (the 3 repeats forever) The part that repeats is usually shown by placing dots over the first and last digits of the repeating pattern, or sometimes a line over the pattern. Also called a "Repeating Decimal".

OpenStudy (unklerhaukus):

the as n approaches infinity the nth term approaches is 4/9, the sum always increases with successive terms